OpenStudy (anonymous):

1 3/34 convert improper fraction or mixed number

OpenStudy (saifoo.khan):

37/34

OpenStudy (anonymous):

crap i mean 1 3/4

OpenStudy (anonymous):

37/34

OpenStudy (lalaly):

\[\frac{(4 \times 1)+3}{4}\]

OpenStudy (lalaly):

=\[\frac{7}{4}\]

OpenStudy (amistre64):